Set up an algebraic equation that could represent the following life situation:

Mpho is trying to make money to help pay for his studies at the
university by selling hot dogs from a hot-dog cart at a soccer
stadium during major games. He pays the cart owner R75 per
night for the use of the cart. He sells hot dogs for R16,50 each.
His costs for the hot dogs, napkins, etc. are R9,60 per hot dog on
average. Calculate the number of hotdogs that Mpho has to sell
to break even (make a zero profit).

Draw the situation for which this equation was set up. You should be able to use this
drawing to explain the situation clearly to a learner as part of remedial work.
1
Expert's answer
2020-06-01T19:15:50-0400

"BE=\frac{FC}{P-VC}"

BE- the break even point

FC- fixed cost

P-price

VC- variable cost

"BE=\frac{75}{16.5-9.6}=10.87=11\ hot\ dogs"



Q1- quantity where incom less than expenses Mpho is losing money

Q0- in this point Mpho doesn't lose or earn money (11)

Q2- here Mpho earns